OpenStudy (anonymous):

How many ways are there to choose a committee of 3 people from a group of 9 people? can someone explain, please?

OpenStudy (anonymous):

This is a part of a branch of mathematics called "Combinatorics" Since order here is not important, we can use 'Combination' to find out the answer \[C(n,r) = \frac{ n! }{ (n-r)!(r)! }\] Where n is the total number of possible options, and r is the number to be chosen. So your answer can be found by :- \[C(9,3) = \frac{ 9! }{ (9-3)! (3)! }\] So, \[C(9,3) = 84\]

OpenStudy (anonymous):

You can find it from this way :!

OpenStudy (anonymous):



OpenStudy (anonymous):

wait how did you get 84 from 9 and 3?

OpenStudy (anonymous):

I hope you are familiar with the factorial notation. An exclamation sign means a factorial. A factorial of number n can be interpreted as \[n \times n-1 \times n-2 \times n-3 ... 1\] So, a factorial of nine would be \[9 \times 8 \times 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1\]

OpenStudy (anonymous):

ohhh!! okay i get it! thank you so much!(: